# PriorDiffusion: Leverage Language Prior in Diffusion Models for Monocular Depth Estimation

### v3

@@ 3 Method
Diffusion formulation. The pipeline of PriorDiffusion is shown in Figure 3. To use a diffusion model for monocular depth estimation, following the formulation of  [11, 29, 57, 58, 83, 30], we adopt the formulation of Denoising Diffusion Probabilistic Model (DDPM) [26] for the forward diffusion process, reverse diffusion process, training objective and inference mentioned below. A latent variable $\mathbf{z}_{t}$ $\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ is defined to represent a noisy version of the depth map in the latent space at diffusion timestep $t$. The ground truch depth map $y^{*}$ is encoded into a latent representation $\mathbf{z_{y}}$ as supervision using a frozen VAE encoder $E$: $\mathbf{z_{y}}=E(y^{*})$. The diffusion latent after the whole denoising process $\mathbf{z_{0}}$ is decoded back to the image space using the frozen VAE decoder $D$: $\hat{y}=D(\mathbf{z}_{0})$.
Figure 5: Visualization on KITTI. PriorDiffusion predicts better depth for described objects, even when parts of the object are almost visible in the image (such as the parking car in the first column). The model is guided for specific described regions, such as a sign at a distance, potentially enhancing the safety of self-driving systems that rely solely on vision sensors.
@@
where $\mu_{\theta}$ and $\Sigma_{\theta}$ is the diffusion model, that predicts the mean and variance conditioned on $\mathbf{z}_{t}$, $t$, the input image $x$ and the corresponding text $\mathbf{c}$. We use the denoising U-Net from Stable Diffusion v2 [54] to initialize is the diffusion model. Specifically, given a text caption $\textbf{c}=\{c_{1},c_{2},...\}$, we first encode it through a frozen CLIP text encoder, then feed it into the diffusion model. Given an image $x$, we encode it using the same VAE encoder $E(x)$ that encoded the depth map, concatenated with the depth latent $\mathbf{z}_{t}$, then feed it into the diffusion model.
Training objective. Our training objective is derived using the reparameterization trick and variational bounds, which involves predicting the noise added at each step. The loss function is:
@@
[…]
Table 1: Quantitative comparison. PriorDiffusion achieves performance on par with other state-of-the-art affine-invariant depth estimators that use significantly more training images, and outperforms methods trained with a similar number of images. Metrics are reported as percentages. * indicates the results of the model that we re-implemented. Note that our method converges faster compared with Marigold, and is even better with only two-thirds iterations, attributing the language prior which provides a stronger constraint for the fast convergence of the diffusion trajectory. For fair comparison, PriorDiffusion and Marigold are evaluated without ensemble.
@@
where $\alpha_{t}$ and $\bar{\alpha}_{t}$ are derived from the noise schedule $\beta_{t}$. This reverse diffusion continues until $t=0$, at which point the fully denoised latent $\mathbf{z}_{0}$ is reached. The predicted depth map $\hat{y}$ is obtained by decoding $\mathbf{z}_{0}$ back into the image space using the pre-trained and frozen VAE decoder:$\hat{y}=D(\mathbf{z}_{0}).$
Figure 6: Controllable depth estimation with language. Each time, we append the newly added sentence to the end of the original sentence and repeat the inference process. Is shows that by providing more details in the language, our method predicts more accurate depth for those specified regions or objects.
@@ 4 Experiments
Training datasets. We adopt the training scheme of Marigold [30], and train our PriorDiffusion on two synthetic datasets that cover both indoor and outdoor scenes. The first dataset utilized is HyperSim [53], a photorealistic compilation of 461 indoor scenes. From HyperSim, approximately 54,000 samples is selected from 365 scenes, adhering to the official data split and excluding any incomplete samples. RGB images and their corresponding depth maps are resized to a resolution of 480 × 640 pixels. The second dataset is Virtual KITTI [3, 18], which comprises synthetic street scenes across five distinct settings, with varied conditions such as different weather patterns and camera perspectives. We specifically use the Virtual KITTI 2 [3] version. For training, we select four scenes, totaling approximately 20,000 samples.
Depth normalization for training. Following [51, 68, 30, 50], to standardize the ground truth depth maps $\mathbf{d}$, we apply a linear normalization ensuring that the depth values primarily fall within the range $[-1,1]$, aligning with the input value range of the VAE. It ensures an affine-invariant depth representation that remains robust to data statistics, guaranteeing that all scenes are constrained by near and far planes with extreme depth values. The affine transformation for normalization is defined as:
$\tilde{y*}=\left(\frac{y^{*}-y_{2}}{y_{98}-y_{2}}-0.5\right)\times 2,$ (1)
where $y_{2}$ and $y_{98}$ represent the 2% and 98% percentiles of the depth maps, respectively.
@@
Evaluation datasets. We assess the performance of our PriorDiffusion using 4 real-world datasets that were not part of its training data, following the configuration as Marigold [30]. Both NYUv2 [60] and ScanNet [9] consist of indoor scenes captured with an RGB-D Kinect sensor. For NYUv2, we use the specified test set, containing 654 images. With ScanNet, we randomly select 800 images from the 312 official validation scenes for evaluation. The KITTI dataset [22, 23] features outdoor street scenes with sparse metric depth captured by a LiDAR sensor, where we apply the Eigen test split [13] consisting of 652 images, following the evaluation protocol of [14]. For ETH3D [59], which are also based on LiDAR measurements, we use all 454 samples that include ground truth depth maps. Following the affine-invariant depth evaluation protocol [75, 51, 68, 30, 50], we evaluate first-order threshold accuracy $\delta_{1}$ and mean absolute relative error $AbsRel$. Details of the metric are provided in the supplementary material.
@@
Obtain language description. To implement our method, we require human-provided text descriptions for each image. Since standard benchmarks do not include such descriptions, we propose using off-the-shelf models to generate text descriptions to simulate those a human would provide. For this purpose, we use the visual question-answering model LLaVA v1.6 [40] to generate 10 text descriptions for each training image using different prompts, and 1 text description for each testing image for evaluation. The details of prompt design are shown in the supplementary material. In each training step, we randomly select 1 text out of 10 that aligns with the input image.
Implementation details. We follow the setup of Marigold [30]. Our method is implemented using PyTorch, using Stable Diffusion v2 [54] as the backbone model. During training, we use the DDPM noise scheduler [26] with 1000 diffusion steps. For inference, we use the DDIM scheduler [61] with 50 sampling steps. Our training process spans 30,000 iterations, with an effective batch size of 32 achieved through gradient accumulation over 16 steps (each with a per-step batch size of 2) to facilitate single-GPU execution. Optimization is performed using the Adam optimizer with a learning rate of $3\cdot 10^{-5}$. Additionally, random horizontal flipping is applied as a data augmentation. Training to convergence, which typically occurs after 20,000 iterations for our method, required approximately four days on an Nvidia RTX 3090 GPU.
Quantitative comparison. Shown in Table 3, PriorDiffusion matches the performance of state-of-the-art affine-invariant depth estimators that require significantly more training images, while outperforming methods trained with a similar number of images. Another significant advantage of PriorDiffusion is its faster convergence. Compared to Marigold, even at 20,000 iterations, PriorDiffusion achieves performance metrics comparable to or better than Marigold’s results at 30,000 iterations. This is also demonstrated in Figure 7. This can be attributed to the language prior integrated into our model, which provides additional semantic and geometric prior serving as an extra constrain that speed up diffusion training. An intuitive explanation is that the model expends less effort in learning the geometric properties and spatial relationships of objects from the image, which is high-dimensional and computationally intensive to process. Instead, much of these features are already embedded in the language input, where these features are represented in a condensed and low-dimensional format, which is easier to learn.
@@
Qualitative comparison. Besides the overall quantitative improvement, we claim that language prior can also guide the model’s attention to better perceive the region that is specified in the language description, which aligns more with users’ interests. The visualization for NYUv2 [60] is shown in Figure 4, where our PriorDiffusion achieves more accurate depth prediction for a given input image compared to Margold, especially for instances highlighted in the language description. For instance, in the first row, the soap dispenser is barely distinguishable in the input image due to its color and texture blending with the background. Vision-based depth estimators like Marigold struggle to identify it, resulting in a depth prediction that matches the background. In contrast, the language description explicitly includes the term “a soap dispenser”, providing the model with a geometric and spatial prior that indicates the existence of a soap dispenser and what it should look like. Consequently, our PriorDiffusion can more accurately perceive the depth of the soap dispenser. Similarly, as shown in Figure 5, we visualize results for the KITTI [22] dataset, where objects such as a parked car or a sign are barely discernible from the visual signal alone. As shown in Figure 6, by incorporating more detailed language descriptions, depth predictions are enhanced for the specified regions or objects.
@@
Faster training convergence. One of the key advantages of PriorDiffusion is its faster convergence compared to Marigold. As shown in Figure 7 here, the improved convergence can be attributed to the integration of a language prior in our model, which provides additional semantic and geometric constraints that help accelerate the diffusion process. Due to computation overhead, following the same configuration of Marigold, we use a subset of KITTI and NYUv2 for training-time evaluation.
@@
Figure 7: Convergence speed comparison. PriorDiffusion converges faster during training compared with Marigold, since it’s more efficient to learn 3D property from condensed, low-dimensional language feature.
@@
Figure 8: Performance under different denoising steps. PriorDiffusion consistently outperforms Marigold across various denoising steps, with a faster convergence.
@@
Fewer denoising steps. As demonstrated in Figure 8, and detailed in Table 4 in the supplementary material, we evaluate PriorDiffusion and Marigold with different denoising steps during inference. PriorDiffusion consistently outperforms Marigold across different denoising steps. Remarkably, PriorDiffusion converges in just 10 steps, while Marigold takes 25 steps to achieve convergence. This indicates that the language prior serves as an effective initialization of the diffusion trajectory, enabling faster convergence.
@@
Better perception in small areas. To demonstrate a better perception for small objects and areas mentioned in text, as shown in Figure 4. We use MaskDINO to obtain panoptic segmentation masks for all images in NYUv2 test set, then evaluate the performance for masks that occupied less than 5%, 10%, and 20% of an image. Table 2 shows that small areas are harder to perceive compared with large areas, and PriorDiffusion obtains better improvement over small areas, as language prior provides geometric and semantic properties that guide the diffusion model to better perceive small areas and tiny structures.
@@
NYUv2 Standard Small Area (5%) Small Area (10%) Small Area (20%) $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 95.7 6.1 91.7 9.0 92.2 8.4 93.9 7.3 Ours 95.9 5.9 92.8 8.3 93.1 7.9 94.6 6.8
@@
Table 2: Better perception in small areas. We evaluate the performance for panoptic segmentation masks that occupied less than 5%, 10%, and 20% of an image. PriorDiffusion obtains better improvement over small areas due to geometric properties of those areas offered by the language prior.
Figure 9: Failure cases. Incorrect text (here we replace ”a bookshelf with glass” with ”a window with curtains”) misguides PriorDiffusion to perceive the geometry of objects.
@@ 5 Discussion
Conclusion. We present PriorDiffusion, a novel approach that integrates language prior with diffusion models to enhance monocular depth estimation. By leveraging geometric prior learned by text-to-image model that associates with language, PriorDiffusion overcomes traditional ambiguities and visual nuisances associated with monocular depth estimation, achieving state-of-the-art performance across various zero-shot benchmarks, while maintaining a faster convergence for training and inference diffusion trajectory.
Limitation. One primary limitation is its dependence on the accuracy and detail of the provided language descriptions. Shown in Figure 9, ambiguous or misleading text inputs from human users can result in suboptimal depth predictions and potentially compromise the model’s performance. Last, our method requires language input from users, which may not always be feasible. In cases where user-provided descriptions are unavailable, as demonstrated in Figure 5 in the supplementary material, our method can still function with standardized template prompts. Another potential solution is to generate language descriptions on the fly using a vision-language model [10], an image captioner [27], or by following methods like RSA [75] to create structured text from the output of visual algorithm [86, 33]. While we have not explored this approach in this paper, it presents a promising avenue for future research.

### v4

@@ 3 Formulation
Diffusion formulation. The pipeline to integrate text is shown in Figure 3. Our way to use diffusion models for monocular depth estimation follows the formulation of  [duan2023diffusiondepth, ji2023ddp, saxena2023monocular, saxena2024surprising, zhao2023unleashing, Marigold], adopting the formulation of Denoising Diffusion Probabilistic Model (DDPM) [DDPM] for the forward diffusion process, reverse diffusion process, training objective and inference mentioned below. A latent variable $\mathbf{z}_{t}$ $\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ is defined to represent a noisy version of the depth map in the latent space at diffusion timestep $t$. The ground truch depth map $y^{*}$ is encoded into a latent representation $\mathbf{z_{y}}$ as supervision using a frozen VAE encoder $E$: $\mathbf{z_{y}}=E(y^{*})$. The diffusion latent after the whole denoising process $\mathbf{z_{0}}$ is decoded back to the image space using the frozen VAE decoder $D$: $\hat{y}=D(\mathbf{z}_{0})$.
Figure 4: Visualization on NYUv2. Compared to the Marigold baseline, integrating language demonstrates more accurate depth prediction for a given input image, particularly for instances specified in the language description (marked in red). This is achieved by providing additional language conditions for the semantic and geometric characteristics of specified objects. It’s particularly beneficial for ambiguous or insignificant areas that are easily neglected by visual signals, like “a soap dispenser” in the first row, and “two black lamps with circular bases” in the second row.
Figure 5: Visualization on KITTI. Integrating language allows predicting better depth for described objects, even when parts of the object are almost invisible in the image (such as the parked car in the first column). Additional semantic and geometrical conditions are provided for the described ambiguous and insignificant regions, such as a sign at a distance, potentially enhancing the safety of self-driving systems that rely solely on vision sensors.
@@
Method Training Images NYUv2 KITTI ETH3D ScanNet DIODE $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ GeoWizard [fu2024geowizard] 280K 96.3 5.6 82.0 14.4 95.8 6.6 95.0 6.4 79.2 29.7 DPT [DPT] 1.4M 90.3 9.8 90.1 10.0 94.6 7.8 93.4 8.2 75.8 18.2 Omnidata [eftekhar2021omnidata] 12.2M 94.5 7.4 83.5 14.9 77.8 16.6 93.6 7.5 74.2 33.9 Depth Anything v2 [DepthAnythingV2] 62.6M 97.9 4.4 94.8 7.5 86.2 13.2 97.8 4.2 95.4 6.5 Depth Anything [DepthAnything] 63.5M 98.1 4.3 94.7 7.6 88.2 12.7 98.1 4.3 95.2 6.6 Marigold [Marigold] 74K 95.9 6.0 90.4 10.5 95.1 7.1 94.5 6.9 77.2 31.0 Marigold 74K 95.7 6.1 89.7 10.7 95.4 6.9 94.0 7.3 77.0 31.3 Marigold + Text (Train Only) 74K 95.5 6.2 89.3 10.9 95.0 6.9 92.6 8.1 77.3 30.7 Marigold + Text (Infer Only) 74K 95.7 6.1 89.8 10.6 95.3 6.9 93.8 7.3 77.1 31.0 Marigold + Text (Train & Infer) 74K 95.9 5.9 90.6 10.4 95.7 6.5 94.9 6.7 78.9 29.8 Lotus-D [he2024lotus] 59K 97.2 5.1 93.1 8.1 97.0 6.1 96.5 5.5 73.8 22.8 Lotus-D 59K 96.6 5.6 92.2 8.7 96.8 6.1 96.0 6.0 74.1 22.2 Lotus-D + Text (Train Only) 59K 96.2 5.8 91.7 8.9 96.9 6.1 95.1 6.4 73.7 24.0 Lotus-D + Text (Infer Only) 59K 96.7 5.6 92.6 8.5 96.7 6.2 96.1 5.8 74.1 22.4 Lotus-D + Text (Train & Infer) 59K 96.8 5.4 93.0 8.4 97.0 6.0 96.6 5.6 74.2 22.0 Lotus-G [he2024lotus] 59K 96.8 5.4 92.2 8.5 97.0 5.9 95.7 5.9 72.9 22.9 Lotus-G 59K 95.2 6.7 92.2 8.9 95.7 9.2 93.7 7.6 71.7 25.4 Lotus-G + Text (Train Only) 59K 95.8 6.3 91.9 9.2 95.6 12.5 93.7 7.3 71.9 25.3 Lotus-G + Text (Infer Only) 59K 95.2 6.7 92.2 9.0 95.5 9.6 94.1 7.3 71.6 28.7 Lotus-G + Text (Train & Infer) 59K 96.3 5.9 92.8 8.6 96.3 9.0 95.3 6.4 72.5 24.3 E2E-FT (Stable Diffusion) [garcia2024fine] 74K 96.5 5.4 92.1 9.6 95.9 6.4 96.5 5.8 77.6 30.3 E2E-FT (Stable Diffusion) 74K 95.4 6.9 90.1 10.5 94.1 8.1 94.6 7.7 76.4 31.0 E2E-FT + Text (Train Only) 74K 95.6 6.9 90.5 10.3 93.3 8.8 95.0 7.5 76.5 31.5 E2E-FT + Text (Infer Only) 74K 96.0 6.4 90.8 10.2 94.2 8.2 95.4 7.1 76.4 31.4 E2E-FT + Text (Train & Infer) 74K 96.3 6.2 91.7 9.7 94.7 7.8 95.0 7.5 77.0 31.4
Table 1: Quantitative comparison. Integrating language generally outperforms the baselines. In some cases, integrating language only during training or only during inference can also lead to performance improvements. Metrics are reported as percentages. * indicates results that we re-trained and re-evaluated using their open-sourced code. Due to computational overhead, all models are evaluated without ensembling. We gray out results that were not reproducible with the released code and models. Unless otherwise specified, in this paper, integrating text means integrating text in both training and inference.
@@
where $\mu_{\theta}$ and $\Sigma_{\theta}$ is the diffusion model, that predicts the mean and variance conditioned on $\mathbf{z}_{t}$, $t$, the input image $x$ and the corresponding text $\mathbf{c}$. The diffusion model is initialized by the denoising U-Net from Stable Diffusion v2 [StableDiffusion]. Specifically, given a text caption $\textbf{c}=\{c_{1},c_{2},$...\}, we first encode it through a frozen CLIP text encoder, then feed it into the diffusion model. Given an image $x$, we encode it using the same VAE encoder $E(x)$ that encoded the depth map, concatenated with the depth latent $\mathbf{z}_{t}$, then feed it into the diffusion model.
Training objective. The training objective is derived using the reparameterization trick and variational bounds, which involves predicting the noise added at each step. The loss function is:
@@
[…]
Unless otherwise specified,“integrating text” in this paper is in both diffusion training and inference.
Figure 6: Iterative depth refinement with language. Each time, we append the newly added sentence to the end of the original sentence and repeat the inference process. Is shows that by providing more details in the language, depth prediction can be iteratively refined, especially for specified regions or objects.
@@ 4 Experiments
Training datasets. The training scheme follows the same ones proposed in Marigold [Marigold], Lotus [he2024lotus], and E2E-FT [garcia2024fine]. All of them are trained on two synthetic datasets that cover both indoor and outdoor scenes. The first dataset utilized is HyperSim [roberts2021hypersim], a photorealistic compilation of 461 indoor scenes. Marigold and E2E-FT approximately select 54,000 samples, and Lotus filters out incomplete samples to keep approximately 39,000 samples. The second dataset is Virtual KITTI [VirtualKITTI, gaidon2016virtual], which comprises synthetic street scenes across five distinct settings, with varied conditions such as different weather patterns and camera perspectives. The Virtual KITTI 2 [VirtualKITTI] version is used, and for all Marigold, Lotus, and E2E-FT, four scenes are selected with totaling approximately 20,000 samples.
@@
Evaluation datasets. Trained models are evaluated using 5 real-world datasets that were not part of their training data: NYUv2 [NYU-Depth-V2], ScanNet [ScanNet], KITTI dataset [KITTI, KITTI_1, Eigen-Split], ETH3D [ETH3D], and DIODE [DIODE], following Marigold [Marigold], Lotus [he2024lotus], and E2E-FT [garcia2024fine]. Following the affine-invariant depth evaluation protocol [zeng2024rsa, DPT, DepthAnything, Marigold, midas], we evaluate first-order threshold accuracy $\delta_{1}$ and mean absolute relative error $AbsRel$. Details of the evaluation data and metric are provided in the Supplementary Material.
@@
Obtain language description. To implement our method, we require human-provided text descriptions for each image. Since standard benchmarks do not include such descriptions, we propose using off-the-shelf models to generate text descriptions to simulate those a human would provide. For this purpose, we use Vision Language Models to generate text. Specifically, we use LLaVA v1.6 [liu2023improved] for Marigold, and InternVL3-8B [zhu2025internvl3] for Lotus and E2E-FT. Each training and testing image is generated with one text descriptions. Details of obtaining language descriptions can be found in the Supplementary Material.
Implementation details. We follow the setup of original Marigold [Marigold], Lotus [he2024lotus], and E2E-FT [garcia2024fine]. Details are provided in the supplementary materials.
Quantitative comparison. Shown in Table 1, after integrating text, the accuracy of diffusion-based monocular depth estimators are generally improved over the baseline. We also find that in some cases, integrating language only during training or only during inference can lead to performance improvements. It shows that language conditions can potentially regularize the diffusion learning process and may also guide depth prediction in a zero-shot manner.
@@
Qualitative comparison. Besides the overall quantitative improvement, we find that language can particularly benefit specified insignificant or ambiguous regions. The visualization for NYUv2 [NYU-Depth-V2] is shown in Figure 4, where integrating text achieves more accurate depth prediction for a given input image compared to the baseline, especially for instances highlighted in the language description. For example, in the first row, the soap dispenser is barely distinguishable in the input image due to its color and texture blending with the background. Vision-based depth estimators like vanilla Marigold struggle to identify it, resulting in a depth prediction that matches the background. In contrast, the language description explicitly includes the term “a soap dispenser,” providing the model with extra conditions that indicate the existence of a soap dispenser and what it should look like, improving its depth estimation. Similarly, as shown in Figure 5, we visualize results for the KITTI [KITTI] dataset, where objects such as a parked car or a sign are barely discernible from the visual signal alone. As shown in Figure 6, by incorporating more detailed language descriptions, depth predictions are enhanced for the specified regions or objects.
@@
Faster training convergence. One of the key advantages of integrating text is its faster convergence compared to Marigold. As shown in Figure 7 here, the improved convergence can be attributed to the integration of language in the diffusion model, which provides additional semantic and geometric constraints that help to learn the diffusion process. Due to computation overhead, following the same configuration of Marigold, we use a subset of KITTI and NYUv2 for training-time evaluation.
@@
Figure 7: Convergence speed comparison. Integrating language converges faster during training compared with the Marigold baseline.
@@
Figure 8: Performance under different denoising steps. Integrating language consistently outperforms the Marigold baseline across various denoising steps, with a faster convergence.
@@
Fewer denoising steps. As demonstrated in Figure 8, and detailed in Table 4 in the supplementary material, we evaluate Marigold with integrating text and the baseline using different denoising steps during inference. We find that integrating text consistently outperforms the baseline across different denoising steps. Remarkably, integrating converges in just 10 steps, while the baseline takes 25 steps to achieve convergence. One explanation is that language provides additional semantic and geometric constraints that speed up the diffusion process.
@@
Better perception in small areas. To demonstrate a better perception for small objects and areas mentioned in text, as shown in Figure 4. We use MaskDINO to obtain panoptic segmentation masks for all images in NYUv2 test set, then evaluate the performance for masks that occupied less than 5%, 10%, and 20% of an image. Table 2 shows that small areas are harder to perceive compared with large areas, and integrating text obtains better improvement over small areas, as language provides a condition for geometric and semantic properties that guide the diffusion model to better perceive small areas and tiny structures.
@@
NYUv2 Standard Small Area (5%) Small Area (10%) Small Area (20%) $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 95.7 6.1 91.7 9.0 92.2 8.4 93.9 7.3 Marigold + Text 95.9 5.9 92.8 8.3 93.1 7.9 94.6 6.8
@@
Table 2: Better perception in small areas. We evaluate the performance for panoptic segmentation masks that occupied less than 5%, 10%, and 20% of an image. Integrating language results in better depth estimation in small areas.
Figure 9: Incorrect text misguides models. After replacing the correct description “a bookshelf with glass” with a wrong description “a window with curtains”, the model failed to perceive the structure of the bookshelf behind the glass.
@@ 5 Discussion
Conclusion. We investigate the influence of integrating language into diffusion-based depth estimators and find that it enhances monocular depth estimation. Our explanation is that language provides an additional condition (learned during the text-to-image pre-training of diffusion models) associated with plausible 3D scenes, thus reducing the solution space for depth estimation.
Limitation. One primary limitation is its dependence on the accuracy and detail of the provided language descriptions. Shown in Figure 9, ambiguous or misleading text inputs from human users can result in suboptimal depth predictions and potentially compromise the model’s performance. Although human-provided descriptions are typically more accurate than automatically generated ones, they can still be incomplete or contain subtle ambiguities. A potential solution is to incorporate a language-robustness module, such as uncertainty estimation or consistency-based filtering, to correct or filter out incorrect descriptions.
[…]